# Quantifying Edits Decay in Fine-tuned LLMs

### v1

@@ 4.2 Qualitative Analysis
We conduct a qualitative analysis by manually examining examples of different model behaviors on target edits after fine-tuning. Table 4 summarizes several identified patterns. First, we observe stable cases where successful edits persist after fine-tuning. These Stable Edits typically involve frequent lexical items as the targets, such as “English”, “Islam”, and “piano”. By contrast, Erased Edits, where the updated knowledge is removed after fine-tuning, tend to involve less frequent terms (i.e., “Lecanorales”), suggesting that frequency and entrenchment of the target knowledge potentially influence the stability of edits. We also observe this pattern in Emergent Edits where unsuccessful edits become successful ones after FT, that the target knowledge involves high-frequency tokens. For instance, when querying the mother tongue of Danielle Darrieux, the expected answer from $M_{\text{ed}}$ is English, but the actual output is “United States”. After fine-tuning, however, $M_{\text{ed\_ft}}$ produces “English”, which is a frequent word.
@@ 4.3 Only Fine-tuning Edited or Non-Edited Layers
As discussed in Section 2.3, our motivation also lies in understanding how to remove potentially harmful edits and how to preserve beneficial ones to avoid repeated editing. At the same time, as shown above, fine-tuning can remove edits from the edited model. Taken together, we propose two hypotheses: (1) FT edited layers can only effectively remove edits; (2) FT non-edited layers can preserve edits.
To examine this, we set two experimental groups: fine-tuning edited layers only and fine-tuning non-edited layers only. For our experiments, we adopt Llama2 and GPT-J as base models, MEMIT and AlphaEdit as KE methods, and LoRA and DoRA as fine-tuning approaches. Specifically, we edit six layers from layer 3 to layer 8 for GPT-J and layers 484{-}8 for Llama2, following Meng et al. (2023) and Wang et al. (2025).
Table 5: KE performance (%) of Llama2 being edited using AlphaEdit on COUNTERFACT dataset, and then being fine-tuned with selective layers. MM1 for model without editing or fine-tuning; $M_{\textit{ed}}$1 for model edited but not fine-tuned; $M_{\textit{ed\_ft\_all}}$3 for edited and then fine-tuned with all layers; $M_{\textit{ed\_ft\_edited}}$4 for edited and then fine-tuned with edited layers; $M_{\textit{ed\_ft\_non-edited}}$5 for edited and then fine-tuned with non-edited layers.
KE performance Llama2 100 Edits 1000 Edits MM1 $M_{e}d$2 $M_{\textit{ed\_ft\_all}}$3 $M_{\textit{ed\_ft\_edited}}$4 $M_{\textit{ed\_ft\_non-edited}}$5 $M_{e}d$ $M_{\textit{ed\_ft\_all}}$ $M_{\textit{ed\_ft\_edited}}$ $M_{\textit{ed\_ft\_non-edited}}$ ES 20.00 96.00 98.00 66.00 72.00 100.00 90.55 57.80 60.30 PS 35.00 87.50 93.00 68.00 64.00 95.75 77.03 60.95 54.40 NS 69.00 76.60 76.60 83.60 82.20 72.44 73.20 79.52 80.31
@@
First, we find that fine-tuning only the edited layers can remove more prior edits than fine-tuning all layers. As illustrated in Table 5, in the case of 100 Edits, between $M_{\text{ed\_ft\_all}}$ and $M_{\text{ed\_ft\_edited}}$, $M_{\text{ed\_ft\_edited}}$ shows a larger drop of editing performance across all three editing metrics (ES, PS, NS). For example, ES of $M_{\text{ed\_ft\_edited}}$ drops to 6666% while ES of $M_{\text{ed\_ft\_all}}$ rises to 9898%, close to 9696% of $M_{\text{ed}}$. However, fine-tuning only edited layers can result in a slight loss of downstream performance. For instance, in the case with 1000 edits in Table 6, performance on BoolQ decreased 3.31% from 71.44% to 68.13%. The only exception is HellaSwag, where performance drops sharply from 89.00% to 32.10%. This underscores the trade-off between effectively removing edits and the risk of losing downstream task performance.
@@
To test whether edits can be preserved by fine-tuning only the non-edited layers, we compare $M_{ed\_ft\_all}$ with $M_{ed\_ft\_non-edited}$ in Table 5. The results are negative: fine-tuning non-edited layers provides no benefit in preserving edits. For example, with 100 edits using AlphaEdit on Llama2, $M_{ed\_ft\_non-edited}$ shows a significant decline in ES from 98% to 72%, whereas $M_{ed\_ft\_all}$ maintains an ES of 96%, close to its pre-FT value. Evaluations on paraphrased prompts yield similar results, with $M_{ed\_ft\_non-edited}$ exhibiting greater degradation.
@@ Discussion
Our experiments yield several interesting findings. First, fine-tuning only the edited layers removes prior edits more effectively, but at the cost of reduced downstream performance in certain scenarios. One possible reason can be that parameters directly involved in edits are highly sensitive to fine-tuning and thus fragile under further adaptation. Second, fine-tuning only the non-edited layers does not preserve edits either, suggesting that edited knowledge is not strictly localized but instead distributed across the network. Taken together, these findings provide a feasible approach for removing unwanted edits by fine-tuning only edited layers. However, this method also highlights the necessity of considering model editing in the whole application pipeline of LLMs, as users may encounter trade-offs between edit removal and downstream performance.
@@ 4.4 Editing performance of Deepseek
In our editing experiments, both AlphaEdit and MEMIT perform poorly on DeepSeek. We verified this result using editing layers identified via causal tracing as well as the default Llama settings, the backbone of the distilled DeepSeek model used here. Detailed results are provided in Appendix G. The limited number of successful edits makes it difficult to fairly assess the impact of fine-tuning on editing, rendering further experiments unnecessary. This underscores the lack of robustness in current KE methods and their unsuitability for emerging models such as DeepSeek.
4.5 KE impact on FT Performance on Downstream
To assess the impact of editing on subsequent fine-tuning, we compare the downstream performance between $M_{\text{ft}}$ and $M_{\text{ed\_ft}}$. We find that KE moderately reduces the effectiveness of subsequent fine-tuning, even when applied beforehand, and the level of impact depends on settings, such as FT methods and KE-related settings. As shown in Table 7, most $M_{\text{ed\_ft}}$ exhibit performance degradation on downstream tasks compared to their counterparts $M_{\text{ft}}$, cross KE methods, and editing datasets. Of the 144 fine-tuned cases, 122 cases experience decline, while the remaining cases exhibit an increase in downstream performance. The largest decline occurs when Llama2 is edited via AlphaEdit on 100 counterfactual facts, followed by LoRA fine-tuning, resulting in a drop of 43.76 p.p. in accuracy rate (from 79.87% to 36.11%).
Further, among the three fine-tuning methods, DoRA is the most affected by KE, as indicated by the predominance of green cells. By contrast, full fine-tuning is the most robust against KE, with 11 out of 36 configurations of even increased performance. An analysis of the relationship between downstream task performance and experiment settings is presented in Appendix H.1.
@@ 5 Conclusion
6 Reproducibility Statement
We provide all necessary resources to facilitate reproducibility of our results. Dataset descriptions and preprocessing steps are detailed in Section 3.1 and Appendix I. Implementation details, model configurations, and training setups are reported in Appendix B and Appendix I. We will release the code to reproduce all experiments once published. Together, these materials ensure that our results can be independently verified.
@@ Appendix A Implementation Details
Appendix B Model configurations
Each column in Table 9 represents a selectable parameter, with experimental settings generated by their Cartesian product. As noted in its caption, configurations with zero edits are consistent across datasets, e.g., Llama2 edited with 0 facts from zsRE and COUNTERFACT dataset are identical. A summary of model details is provided in Table 8. Additionally, DeepSeek is not fine-tuned, with the rationale detailed in Section 4.4.
@@ Appendix C Results validation
As shown in Table 10, the results obtained from our validation experiments closely align with the original values reported in the paper, indicating strong reproducibility and correctness of our re-implementation.
@@
As shown in Figure 3 and Figure 4, validating results are close to original results, indicating the reliability of our outputs.
@@ Appendix D Knowledge Editing metrics
We construct our evaluation using metrics defined in previous works(Meng et al., 2022; Yao et al., 2023). For below equations, P for probability of model G generating certain outputs, $o_{i}$ for answers produced, and $\mathbb{E}$ represents mathematical expectation.
@@
Efficacy measures the proportion of successful edits, defined by whether the model assigns a higher probability to the edited target than to the original answer for a given fact and its corresponding query prompt ($P\text{(edited target))}>P\text{(original answer)}$).
Formulas 2 and 3 present the mathematical definition of Efficacy for the zsRE and COUNTERFACT datasets, respectively.
$\mathbb{E}_{i}\left[o_{i}=\arg\max_{x_{E}}\mathbb{P}_{G}\left(x_{E}\mid p(s_{i},r_{i})\right)\right]$ (2)
$\mathbb{E}_{i}\!\left[\mathbb{P}_{G}\!\left(o_{i}\mid p(s_{i},r_{i})\right)\right]>\mathbb{P}_{G}\!\left(o_{i}^{c}\mid p(s_{i},r_{i})\right)$ (3)
@@
Paraphrase evaluates model’s generalization ability after editing facts. Similar to Efficacy, it is measured by checking whether $P\text{(edited target))}>P\text{(original answer))}$ but under a paraphrased query prompt.
Formulas 4 and 5 present the mathematical definition of Paraphrase for the zsRE and COUNTERFACT datasets, respectively.
$\mathbb{E}_{i}\left[\mathbb{E}_{(s_{i},r_{i})\in\text{paraphrases}(s_{i},r_{i})}\left[o_{i}=\underset{x_{E}}{\arg\max}\,\mathbb{P}_{G}(x_{E}\mid p)\right]\right]$ (4)
$\mathbb{E}_{i}\left[\mathbb{E}_{(s_{i},r_{i})\in\text{paraphrases}(s_{i},r_{i})}\left[\mathbb{P}_{G}[o_{i}\mid p]>\mathbb{P}_{G}[o_{i}^{c}\mid p]\right]\right]$ (5)
@@
Specificity assesses the locality of a knowledge edit by measuring its unwanted impact on facts nearby the facts involved in KE. NS calculates the percentage of unedited facts nearby which remain unchanged after editing.
Formulas 6 and 7 present the mathematical definition of Specificity for the zsRE and COUNTERFACT datasets, respectively.
$\mathbb{E}_{i}\left[\mathbb{E}_{p\in\text{neighborhood prompts}(s_{i};r_{i})}\left[o_{i}^{c}=\underset{x_{E}}{\arg\max}\,\mathbb{P}_{G}\left(x_{E}\mid p\right)\right]\right]$ (6)
$\mathbb{E}_{j}\left[\mathbb{E}_{\text{neighborhood prompts}(s_{i},r_{j})}\left[\mathbb{P}_{G}(o_{i}^{j}\mid p)<\mathbb{P}_{G}(o_{i}^{j}\mid p^{\prime})\right]\right]$ (7)

### v3

@@ 4.2 Qualitative Analysis
We conduct a qualitative analysis by manually examining examples of different model behaviors on target edits after fine-tuning. Tab. 4 summarizes several identified patterns. First, we observe stable cases where successful edits persist after fine-tuning. These Stable Edits typically involve frequent lexical items as the targets, such as “English”, “Islam”, and “piano”. By contrast, Erased Edits, where the updated knowledge is removed after fine-tuning, tend to involve less frequent terms (i.e., “Lecanorales”), suggesting that frequency and entrenchment of the target knowledge potentially influence the stability of edits. We also observe this pattern in Emergent Edits where unsuccessful edits become successful ones after FT, that the target knowledge involves high-frequency tokens. For instance, when querying the mother tongue of Danielle Darrieux, the expected answer from $M_{\text{ed}}$ is English, but the actual output is “United States”. After fine-tuning, however, $M_{\text{ed\_ft}}$ produces “English”, which is a frequent word.
@@ 4.3 Only Fine-tuning Edited or Non-Edited Layers
As discussed in Sec. 2.3, our motivation also lies in understanding how to remove potentially harmful edits and how to preserve beneficial ones to avoid repeated editing. At the same time, as shown above, fine-tuning can remove edits from the edited model. Taken together, we propose two hypotheses: (1) FT edited layers can only effectively remove edits; (2) FT non-edited layers can preserve edits.
To examine this, we set two experimental groups: fine-tuning only the edited layers and fine-tuning only the non-edited layers. For our experiments, we adopt Llama2 and GPT-J as base models, MEMIT and AlphaEdit as KE methods, and LoRA and DoRA as fine-tuning approaches. Specifically, we edit layer 383{-}8 for GPT-J and layers 484{-}8 for Llama2, following Meng et al. (2023) and Wang et al. (2025).
KE performance Llama2 100 Edits 1000 Edits MM1 $M_{\text{ed}}$2 $M_{\textit{ed\_ft\_all}}$3 $M_{\textit{ed\_ft\_edited}}$4 $M_{\textit{ed\_ft\_non-edited}}$5 $M_{\text{ed}}$ $M_{\textit{ed\_ft\_all}}$ $M_{\textit{ed\_ft\_edited}}$ $M_{\textit{ed\_ft\_non-edited}}$ ES 20.00 96.00 98.00 66.00 72.00 100.00 90.55 57.80 60.30 PS 35.00 87.50 93.00 68.00 64.00 95.75 77.03 60.95 54.40 NS 69.00 76.60 76.60 83.60 82.20 72.44 73.20 79.52 80.31 DS 1.77 2.15 81.7 65.43 80.61 4.79 81.00 65.35 72.46
Table 5: KE performance (%) of Llama2 being edited using AlphaEdit on COUNTERFACT dataset, and then being fine-tuned with selective layers. MM1 for model without editing or fine-tuning; $M_{\textit{ed}}$1 for edited-only model; $M_{\textit{ed\_ft\_all}}$3 for edited-then-finetuned with all layers; $M_{\textit{ed\_ft\_edited}}$4 for edited-then-finetuned with edited layers; $M_{\textit{ed\_ft\_non-edited}}$5 for edited-then-finetuned with non-edited layers. ES, NS and PS are KE metrics, DS is the average score of downstream tasks.
@@
First, we find that fine-tuning only the edited layers can remove more prior edits than fine-tuning all layers. As illustrated in Tab. 5, in the case of 100 Edits, between $M_{\text{ed\_ft\_all}}$ and $M_{\text{ed\_ft\_edited}}$, $M_{\text{ed\_ft\_edited}}$ shows a larger drop of editing performance across all three editing metrics (ES, PS, NS). For example, ES of $M_{\text{ed\_ft\_edited}}$ drops to 6666% while ES of $M_{\text{ed\_ft\_all}}$ rises to 9898%, close to 9696% of $M_{\text{ed}}$. However, fine-tuning only edited layers can result in a loss of downstream performance. For instance, in the case with 1000 edits in Tab. 16, performance on BoolQ decreased 3.31% from 71.44% to 68.13%. The only exception is HellaSwag, where performance drops sharply from 89.00% to 32.10%. When jointly considering ES and the overall downstream performance ($\Delta ES$ vs. $\Delta DS$), we observe that although $\Delta ES$ decreases by nearly 45% (from 96% to 66%), the average downstream score of the 100-edit model declines by only 24% (from 81.7% to 65.43%). This underscores the trade-off between effectively removing edits and the risk of losing downstream task performance. If overall downstream performance is not a priority, fine-tuning only the edited layers is an effective strategy for removing unwanted edits.
@@
To test whether edits can be preserved by fine-tuning only the non-edited layers, we compare $M_{ed\_ft\_all}$ with $M_{ed\_ft\_non-edited}$ in Tab. 5. The results are negative: fine-tuning non-edited layers provides no benefit in preserving edits. For example, with 100 edits using AlphaEdit on Llama2, $M_{ed\_ft\_non-edited}$ shows a significant decline in ES from 98% to 72%, whereas $M_{ed\_ft\_all}$ maintains an ES of 96%, close to its pre-FT value. Evaluations on paraphrased prompts yield similar results, with $M_{ed\_ft\_non-edited}$ exhibiting greater degradation. We further investigate whether fine-tuning only the non-edited layers can effectively remove edits. As shown in Tab. 18, this approach preserves stronger downstream performance (80.61 vs. 65.43; 72.46 vs. 65.35, all in %) but erases fewer edits than fine-tuning only the edited layers. These results suggest that fine-tuning non-edited layers can be a supplementary edit-removal strategy.
@@ Discussion
Through above experiments, we have findings as: (i) in edit‑removal, $M_{ed\_ft\_edited}>M_{ed\_ft\_non-edited}>M_{ed\_ft\_all}$; (ii) regarding the effectiveness of fine-tuning, we have $M_{ed\_ft\_all}>M_{ed\_ft\_non-edited}>M_{ed\_ft\_edited}$. Above observation aligns with the distributed representation hypothesis, which posits that factual associations in LLMs emerge from coordinated patterns across many MLP and attention layers (Geva et al., 2023; Dar et al., 2023). Editing, which modifies only a subset of weights, leads to incomplete shifts within these distributed circuits. Fine-tuning can readily disrupt the coordinated structure that supports the edits.
@@ 4.4 Editing performance of Deepseek
In our editing experiments, both AlphaEdit and MEMIT perform poorly on DeepSeek. We verified this result using editing layers identified via causal tracing as well as the default Llama settings, the backbone of the distilled DeepSeek model used here. Detailed results are provided in App. G. The limited number of successful edits makes it difficult to fairly assess the impact of fine-tuning on editing, rendering further experiments unnecessary. This underscores the lack of robustness in current KE methods and their unsuitability for emerging models such as DeepSeek.
4.5 Ablation Analysis
KE impact on FT Performance on Downstream
To assess the impact of editing on subsequent fine-tuning, we compare the downstream performance between $M_{\text{ft}}$ and $M_{\text{ed\_ft}}$. We find that KE moderately reduces the effectiveness of subsequent fine-tuning, even when applied beforehand, and the level of impact depends on settings, such as FT methods and KE-related settings. As shown in Tab. 6, most $M_{\text{ed\_ft}}$ exhibit performance degradation on downstream tasks compared to their counterparts $M_{\text{ft}}$, cross KE methods, and editing datasets. Of the 144 fine-tuned cases, 122 cases experience decline, while the remaining cases exhibit an increase in downstream performance. The largest decline occurs when Llama2 is edited via AlphaEdit on 100 counterfactual facts, followed by LoRA fine-tuning, resulting in a drop of 43.76 p.p. in accuracy rate (from 79.87% to 36.11%). Among the three fine-tuning methods, DoRA is the most severely affected by KE, as indicated by the predominance of green cells, whereas full fine-tuning is the most robust, with 11 out of 36 configurations showing even improved performance. A detailed analysis of downstream task performance across experiment settings is provided in App. H.1.
Catastrophic Forgetting
We further examine whether catastrophic forgetting, rather than KE, drives the observed edit decay. To do so, we compare the edited-then-finetuned model $M_{\text{ed\_ft}}$ with its fine-tuned-only counterpart $M_{\text{ft}}$ on downstream tasks. The results indicate that catastrophic forgetting is unlikely the cause: if it were, both $M_{\text{ed\_ft}}$ and $M_{\text{ft}}$ would show performance drops. Instead, as shown in Tab. 6, $M_{\text{ed\_ft}}$ ($77.95$) achieves comparable downstream performance to $M_{\text{ft}}$ ($1.77$), both outperforming $M_{\text{ed}}$ ($0.83$) and $M$ ($0.83$). This pattern suggests that $M_{\text{ed\_ft}}$ does not have catastrophic forgetting issue.
@@ 4.5 Ablation Analysis
5 Why Editing is Fragile to Fine-tuning
Building on Sec. 4.3, which shows that fine-tuning non-edited layers can degrade editing performance, we note that factual associations in LLMs are encoded via distributed mechanisms across multiple layers and directions within the residual stream (Dar et al., 2023; Geva et al., 2023; Choe et al., 2025). Motivated by this, we investigate whether a knowledge edit induces a coherent shift in activation space and how subsequent fine-tuning affects it. For a model $M$, its edited version $M_{\text{ed}}$, and its edited-then-fine-tuned version $M_{\text{ed\_ft}}$, we analyze activations $h_{\ell}(x)$ at each layer $\ell$ using prompt $x$ from a diagnostic prompt set $\mathcal{X}$, which comprises two groups: prompts that explicitly query the edited knowledge (from the editing dataset, Sec. 3.1) and prompts that do not directly invoke the edited fact (from downstream tasks, Sec. 3.1).
Layer-wise drift
For each prompt $x$, we compute the magnitude of activation changes introduced by editing ($\Delta^{\text{ed}}_{\ell}(x)$), by fine-tuning ($\Delta^{\text{ft}}_{\ell}(x)$) and by both ($\Delta^{\text{ed\_ft}}_{\ell}(x)$), where :
$\Delta^{\text{ed}}_{\ell}(x)$ $=\|h_{\ell}^{M_{\text{ed}}}(x)-h_{\ell}(x)\|_{2}$ (2)
$\Delta^{\text{ft}}_{\ell}(x)$ $=\|h_{\ell}^{M_{\text{ft}}}(x)-h_{\ell}(x)\|_{2}$
$\Delta^{\text{ed\_ft}}_{\ell}(x)$ $=\|h_{\ell}^{M_{\text{ed\_ft}}}(x)-h_{\ell}^{M_{\text{ed}}}(x)\|_{2}$
Figure 3: In layer-wise activation drifts (upper three) for GPT2-XL, GPT-J and Llama2, 3 categories for each model: $M_{\text{ed}}$, $M_{\text{ft}}$ and $M_{\text{ed\_ft}}$. In directional similarities (bottom three), 3 pairs of categories tested for each model: $M_{\text{ed}}$ - $M_{\text{ft}}$, $M_{\text{ed\_ft}}$ - $M_{\text{ft}}$ and $M_{\text{ed\_ft}}$ - $M_{\text{ed}}$. Within the red vertical dash lines are the range of layers being edited. Result specifications in App. I
We compute the arithmetic mean over all prompts and visualize activation changes across layers in Fig. 3. We observe that (i) fine-tuned models ($M_{\text{ft}}$, $M_{\text{ed\_ft}}$) exhibit larger activation changes than non-fine-tuned models, and (ii) edited-only models show changes primarily from the edited layer, while fine-tuning affects a broader range of layers. These findings suggest that edits induce shallow, localized activation perturbations, which can be overwritten by the broader effects of fine-tuning, as discussed in Sec. 4.3.
Directional similarity
To further characterise how fine-tuning interacts with the directions introduced by editing, we compute the cosine similarity between the editing direction and the fine-tuning direction in activation space. We first define the layer-wise displacement vectors using Equ. 6. We then compute the layer-wise directional similarity by averaging similarities for single prompts $x$ ($x\in\mathcal{X}$) at layer $\ell$, as shown in Equ. 6 ($\varepsilon\!=\!10^{-8}$ is used to prevent division by zero):
$\Delta^{\text{ed}}_{\ell}(x)=h_{\ell}^{M_{\text{ed}}}(x)-h_{\ell}(x),$ (3)
$\Delta^{\text{ft}}_{\ell}(x)=h_{\ell}^{M_{\text{ft}}}(x)-h_{\ell}(x),$ (4)
$\Delta^{\text{ed\_ft}}_{\ell}(x)=h_{\ell}^{M_{\text{ed\_ft}}}(x)-h_{\ell}^{M_{\text{ed}}}(x).$ (5)
$\text{sim}_{\ell}(x)=\frac{1}{|\mathcal{X}|}\sum_{x\in\mathcal{X}}\frac{\langle\Delta^{\text{1}}_{\ell}(x),\;\Delta^{\text{2}}_{\ell}(x)\rangle}{\|\Delta^{M_{\text{1}}}_{\ell}(x)\|_{2}\;\|\Delta^{M_{\text{2}}}_{\ell}(x)\|_{2}+\varepsilon}.$ (6)
A value $\text{sim}_{\ell}\!\approx\!1$ would indicate that fine-tuning pushes activations further in the direction of the KE, whereas a negative value suggests they are in the opposite direction. As shown in the bottom row of Fig. 3, fine-tuned models ($M_{\text{ed\_ft}}$ - $M_{\text{ft}}$) shares the lowest similarity, indicating that fine-tuning moves activations in the positive directions nearly orthogonal to the editing direction. This orthogonality helps explain why edits are overwritten even when kept in their original layers.
Discussion
The results show that $M_{\text{ft}}$ and $M_{\text{ed\_ft}}$ exhibit both the largest activation-magnitude changes and the highest activation similarity, indicating that fine-tuning, not KE, overwhelmingly dominates models’ representations. In contrast, $M_{\text{ed}}$ has small, dispersed activation shifts which are deviant to the directions introduced by fine-tuning. The above findings indicates that edited knowledge may be overwritten by fine-tuned knowledge during representation.
@@ Appendix A Implementation Details
Appendix B Model configurations
Each column in Tab. 8 represents a selectable parameter, with experimental settings generated by their Cartesian product. As noted in its caption, configurations with zero edits are consistent across datasets, e.g., Llama2 edited with 0 facts from zsRE and COUNTERFACT dataset are identical. A summary of model details is provided in Tab. 7. Additionally, DeepSeek is not fine-tuned, with the rationale detailed in Sec. 4.4.
@@ Appendix C Results validation
As shown in Tab. 9, the results obtained from our validation experiments closely align with the original values reported in the paper, indicating strong reproducibility and correctness of our re-implementation.
@@
As shown in Fig. 4 and Fig. 5, validating results are close to original results, indicating the reliability of our outputs.
@@ Appendix D Knowledge Editing metrics
We construct our evaluation using metrics defined in previous works(Meng et al., 2022; Yao et al., 2023). For each edit instance $i$ from zsRE in an edited model $M_{\text{ed}}$, we set $s_{i}$ to be the subject, $r_{i}$ to be the relation, and $o_{i}$ to be the target object. We write $p(s_{i},r_{i})$ for the base prompt constructed from $(s_{i},r_{i})$. For COUNTERFACT, we additionally denote by $o_{i}^{c}$ the original (counterpart) object describing the real-world fact, and by $\text{paraphrases}(s_{i},r_{i})$ and $\text{neighborhood}(s_{i},r_{i})$ the sets of paraphrased and neighborhood prompts, respectively. Given a prompt $p$, we use $Pr_{\text{$M_{\text{ed}}$}}(x\mid p)$ for the model’s predicted probability of token $x$.
@@
Efficacy measures the proportion of successful edits. Formulas 7 and 8 determine the successful editing of an editing instance from zsRE and COUNTERFACT datasets, respectively. For zsRE, an edit is considered successful on instance $i$ if the edited model assigns the highest probability to the desired answer $o_{i}$ under the base prompt $p(s_{i},r_{i})$ (Equ. 7). For COUNTERFACT, each edit $i$ specifies a counterfactual object $o_{i}$ to be written and a corresponding real-world object $o_{i}^{c}$. An edit $i$ is considered successful, under the base prompt $p(s_{i},r_{i})$, if the edited model assigns higher probability to the desired counterfactual $o_{i}$ than to the original object $o_{i}^{c}$ (Equ. 8):
$\text{ES}_{i}^{\text{zsRE}}=\mathbf{1}\Bigl(o_{i}=\arg\max_{x}Pr_{\text{$M_{\text{ed}}$}}\!\bigl(x\mid p(s_{i},r_{i})\bigr)\Bigr)$ (7)
$\text{ES}_{i}^{\text{CF}}=\mathbf{1}\Bigl(Pr_{\text{$M_{\text{ed}}$}}\!\bigl(o_{i}\mid p(s_{i},r_{i})\bigr)>Pr_{\text{$M_{\text{ed}}$}}\!\bigl(o_{i}^{c}\mid p(s_{i},r_{i})\bigr)\Bigr)$ (8)
Thus, the overall ES can be calculated as:
$\mathrm{ES}^{\text{CF/zsRE}}=\frac{1}{N}\sum_{i=1}^{N}\text{ES}_{i}^{\text{CF/zsRE}}$ (9)
@@
Paraphrase evaluates model’s generalization ability after editing facts. We consider for each instance $i$ a set of paraphrases $paraphrases(s_{i},r_{i})$ of the base prompt. For zsRE, we evaluate average top-1 accuracy on rephrased prompts $N(s_{i},r_{i})$. For COUNTERFACT, On a rephrased prompt $p\in paraphrases(s_{i},r_{i})$, we declare success if the model again prefers the counterfactual object over the original. Formulas 10 and 11 present the mathematical definition of Paraphrase for the zsRE and COUNTERFACT datasets, respectively:
$\mathrm{PS}^{\text{zsRE}}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}\!\left(o_{i}=\arg\max_{o}Pr_{\text{$M_{\text{ed}}$}}\!\left(o\mid N(s_{i},r_{i})\right)\right)$ (10)
$\text{PS}_{i}^{\text{CF}}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}\Bigl(Pr_{\text{$M_{\text{ed}}$}}\!\bigl(o_{i}\mid p\bigr)>Pr_{\text{$M_{\text{ed}}$}}\!\bigl(o_{i}^{c}\mid p\bigr)\Bigr)$ (11)
@@
Specificity assesses the locality of a knowledge edit by measuring its unwanted impact on facts unrelated to the facts involved in KE. To obtain NS, we consider for each instance $i$ a set of neighborhood prompts $\text{neighborhood}(s_{i},r_{i})$ that should not be affected by the edit. Formulas 12 and 13 present the mathematical definition of Specificity for the zsRE and COUNTERFACT datasets, respectively. For zsRE, $O(s_{i},r_{i})$ represents the unrelated facts:
$\mathrm{NS}^{\text{zsRE}}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}\!\left(o_{i}=\arg\max_{o}Pr_{\text{$M_{\text{ed}}$}}\!\left(o\mid O(s_{i},r_{i})\right)\right)$ (12)
$\text{NS}^{\text{CF}}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}\Bigl(Pr_{\text{$M_{\text{ed}}$}}\!\bigl(o_{i}\mid p\bigr)<Pr_{\text{$M_{\text{ed}}$}}\!\bigl(o_{i}^{c}\mid p\bigr)\Bigr)$ (13)